# SteerDiff: Steering towards Safe Text-to-Image Diffusion Models

### v1

@@ 3.3 Steering Toward Safe Content
To steer the inappropriate prompts toward generating safe images, we propose to learn a linear transformation to the embedding of identified unsafe prompts. This transformation shifts unsafe concepts toward safe ones while preserving original semantic meaning. The intuition behind this approach is that linear transformations of word embeddings can effectively steer the generation style of language models (Han et al., 2023). Since SteerDiff operates in word embedding space, applying these transformations enables us to adjust unsafe embeddings and steer the diffusion model toward producing safe outputs. Specifically, let $E$ denote the embedding of a word, we define the transformation as follows:
${E}_{\text{steered}}=\epsilon W{E}_{\text{unsafe}}+(1-\epsilon){E}_{\text{unsafe}}$ (3)
In this equation, ${E}_{\text{steered}}$ represents the adjusted embedding after steering, ${E}_{\text{unsafe}}$ refers to the original embedding of the unsafe phrases in $T_{\text{unsafe}}$, $\epsilon$ is a steering parameter that controls the intensity of the transformation, and $W$ is a linear transformation matrix learned during training.
The key concept is to map the embedding of inappropriate content to safe content through a linear transformation. By applying Equation 3, we shift the embedding within the semantic space, guiding it toward regions associated with safe imagery. The steering parameter $\epsilon$ offers fine-grained control over the degree of transformation, allowing flexibility in balancing semantic preservation and safety. To learn the transformation matrix $W$, we employ a supervised learning method using a paired dataset of unsafe phrases and their corresponding safe phrases, as described in subsection 3.1. The training process minimizes the following loss function:
$\mathcal{L}=\left|{E}_{\text{safe}}-W\cdot{E}_{\text{unsafe}}\right|^{2}$ (4)
where ${E}_{\text{safe}}$ represents the embedding of the safe phrases in $T_{\text{safe}}$.
3.4 Inference
@@
Once SteerDiff is trained, it can be seamlessly integrated into diffusion models without additional fine-tuning. In practical applications, SteerDiff can effectively detect the presence of blacklisted concepts and steer inappropriate prompts toward generating safe images.
Figure 3: Illustration of the SteerDiff process: A problematic embedding (red arrow) is steered (blue arrow) towards a safe embedding (green arrow) to mitigate the generation of inappropriate content.
@@
Consider a T2I model equipped with a text encoder. As illustrated in Figure 3, a user provides a prompt $p$, which can be either safe or unsafe. The input prompt is first embedded by a text encoder. We define a concept blacklist $C_{\text{unsafe}}$ that contains potentially inappropriate concepts. Then the identifier detects whether prompt $p$ contains any inappropriate concepts $c^{\prime}\in C_{\text{unsafe}}$. If such a concept is detected, a linear transformation is applied to the prompt’s embedding to steer it toward safer content. Specifically, the transformation described in Equation 3 is applied to the embedding of $p$, altering the latent representation to avoid generating inappropriate content. This transformation adjusts the embedding space without compromising the semantics of the original prompt. Once the transformation is applied, the modified embedding is passed through the diffusion model to generate a safe image. The transformed embedding ensures that the resulting image adheres to SFW guidelines while preserving the user’s original intent.
@@
This approach allows SteerDiff serving as an intermediary between the user’s input and the diffusion model, ensuring that the generated images align with ethical and safety standards. Additionally, this steering process is efficient, eliminating the need for model retraining or introducing significant delays in the image generation pipeline. The transformation can also be dynamically adapted as blacklists or safety criteria evolve, making the system robust and scalable for real-world applications.
4 Experiment
4.1 Experimental Setup and Metrics
@@
To evaluate the reduction of inappropriate content, we used Stable Diffusion (SD) v1.4 as the base model to generate three images for each prompt. We selected a general concept set $C_{\text{unsafe}}$, applicable across all categories of Inappropriate Image Prompts (I2P) dataset (Schramowski et al., 2023), as outlined in Equation 1. We evaluated inappropriate image content by combining two classifiers. First, Q16 classifier (Schramowski et al., 2022), which is also used for annotating LAION-5B dataset (Schuhmann et al., 2022), was utilized to detect a broad range of inappropriate content. Second, we applied NudeNet (Praneeth, 2019) to identify sexually explicit content. For nude content, we classified exposed buttocks, exposed breasts, exposed genitalia, and exposed anus content as inappropriate, while allowing otherwise provocative images (Schramowski et al., 2023; Wu et al., 2024; Ma et al., 2024). Unless otherwise specified, an image was classified as inappropriate if at least one of the classifiers flagged it as containing such content.
For the following RQ1, we evaluated SteerDiff on the I2P dataset and compared the probability of generating inappropriate images with state-of-the-art defense model Safe Latent Diffusion (SLD) under its MAX and STRONG configurations. In addition, we compared the probability of generating nude images with Erased Stable Diffusion (ESD) (Gandikota et al., 2023). For RQ2, we selected two state-of-the-art red teaming frameworks: Prompting4Debugging (P4D), representing a white-box attack, and SneakyPrompt, representing a black-box attack. Following standard instructions from P4D and Sneakyprompt, we conducted our evaluation on the I2P dataset using P4D testbed and NSFW_200 dataset (Yang et al., 2024) with Sneakyprompt, utilizing standard configurations provided by developers. For the diffusion process, we used the default configuration as outlined by stable-diffusion (Rombach et al., 2021). This approach ensures our experiments align with recommended practices to maintain consistency across all tests. For RQ3, we assessed image fidelity and text alignment across all models using prompts from the COCO 30K dataset.
4.2 RQ1: How effective is SteerDiff in mitigating generation of inappropriate content?
@@
To investigate the ability of SteerDiff in identifying and steering inappropriate concepts, we started by demonstrating its effectiveness in reducing the generation of explicit content. We compared SteerDiff against ESD, SLD STRONG, and SLD MAX. Next, we expanded the scope of inappropriate concepts to $C_{\text{unsafe}}$ to investigate whether SteerDiff could effectively identify and steer prompts containing a wider range of inappropriate content toward generating safe images. In this evaluation, we evaluated SteerDiff with Stable Diffusion (SD) v1.4, naive blacklist, SLD STRONG, and SLD MAX approach. To minimize randomness and ensure more reliable results, all evaluated methods generated three images per prompt, as generating only one image might coincidentally omit inappropriate content. Notably, we used inappropriate terms from SteerDiff’s training set as the naive blacklist.
4.2.1 Explicit Content Removal
Figure 4: Our method effectively removes sexual content from SD v1.4 on the I2P dataset, outperforming defend methods ESD, SLD STRONG, and SLD MAX. Illustrating the percentage reduced in nudity-classified samples compared to the original SD v1.4 model.
We evaluated the performance of SteerDiff and competitors on the I2P dataset, focusing on the sexual content category. In Figure 4, we compared the percentage change in nudity-classified samples with respect to SD v1.4. Our results show that, across all classes, our method demonstrates a more significant reduction in the generation of explicit content. Specifically, SD v1.4 generated 637 images with exposed body parts on test prompts, SteerDiff, ESD, SLD STRONG, and SLD MAX reduced this number to 104, 389, 208, and 134, respectively.
4.2.2 Inappropriate Content Removal
@@
Table 1: SteerDiff demonstrates the best performance in reducing the probability of generating inappropriate content (where lower values are better). The probabilities shown represent the likelihood of generating images classified as inappropriate by combining the Q16 and NudeNet classifiers across various I2P categories. The best performances are bolded, and the second-best performances are underscored.
@@
SD1.4 27.27 19.05 27.65 30.34 46.29 35.98 18.16 30.17
Blacklist 19.48 14.68 17.99 18.23 21.59 22.31 10.87 17.86
SteerDiff (Ours) 5.63 4.25 2.91 4.74 2.36 6.78 3.99 4.51
ESD - - - - 10.56 - - -
SLD STRONG 6.49 6.80 5.42 5.24 12.67 11.33 3.03 7.97
SLD MAX 3.90 3.76 5.29 2.25 8.38 6.31 2.75 5.15
@@
We further investigated a more comprehensive inappropriate set $C_{\text{unsafe}}$ defined in Equation 1. We began our evaluation by demonstrating the inappropriate generation of SD v1.4 without any safety measures, as well as a basic blacklist-based approach for prompt matching. Table 1 presents the probability of generating inappropriate content for each category. Varying from different categories, SD v1.4 generated inappropriate content with probabilities ranging from 18.16% to 46.29%. The naive blacklist approach slightly reduced probability, but inappropriate content was still generated in 17.86% of cases across all categories. While naive blacklists may mitigate some inappropriate image generation, it remains largely impractical as a comprehensive defense mechanism due to its inability to capture the diverse and evolving nature of unsafe content.
@@
Next, we demonstrated the probability of generating inappropriate content using SteerDiff, SLD STRONG, and SLD MAX across different I2P concepts. As shown in Table $1$ESD data was sourced from Ma et al. (2024)., both SteerDiff and SLD MAX demonstrate the strongest performance, ranking first and second, respectively. Specifically, SteerDiff reduced the probability of generating inappropriate content by over 85%. In particular, SteerDiff outperformed its closest competitor, SLD MAX, in categories of violence, sexual content, and overall inappropriate content. As a result, only 5% of images generated by SteerDiff were still classified as inappropriate. However, it is worth noting that the Q16 and Nudenet classifiers tend to flag images as inappropriate even when problematic content has been significantly reduced. In summary, SteerDiff effectively mitigates the generation of inappropriate content in SD by identifying and modifying unsafe concepts within text embeddings.
4.3 RQ2: Can SteerDiff defend against red-teaming attacks?
While SteerDiff was the state-of-the-art model when evaluating the I2P dataset, the robustness against red-teaming attacks is also a necessary factor of successful defense. Therefore, we committed to further investigating its efficacy towards red-teaming attacks. To this end, we selected state-of-the-art white-box and black-box red-teaming frameworks Prompt4dubugging (P4D) and SneakyPrompt (Chin et al., 2023; Yang et al., 2024), respectively.
Table 2: Performance of various defense methods under no attack, P4D (Chin et al., 2023), and SneakyPrompt (Yang et al., 2024), evaluated using attack success rate. Bold values indicate the highest performance, while underlined values represent the second-highest performance.
@@ 4.3 RQ2: Can SteerDiff defend against red-teaming attacks?
No attack (Nude) No attack (all) white-box black-box
P4D (nude) P4D (all) SneakyPrompt
SteerDiff (Ours) 2.36 4.51 25.36 29.16 7.50
ESD 10.56 - 55.40 - 51.00
SLD STRONG 12.67 7.97 48.21 - 14.50
SLD MAX 8.38 5.15 37.25 30.95 8.50
To investigate the effectiveness of concept removal approaches, we first focused on the “nudity” category, as it is commonly recognized as explicitly harmful in the context of generative models. Specifically, we inspected all safe T2I models for the nudity category in the I2P dataset. As shown in LABEL:T:exp2, both the SLD and ESD exhibited poor performance when countering nudity-related attacks. In particular, over 50% of attacks launched by the P4D method successfully bypassed the ESD, while 48.21% and 37.25% circumvented SLD STRONG and MAX defenses, respectively. In contrast, only 25.36% of attacks successfully bypassed SteerDiff, marking it as the most effective defense against red-teaming attacks targeting nudity. A broader range of image results can be found in Figure 10.
Next, we evaluated the robustness of SLD MAX and SteerDiff across all categories within the I2P dataset under the P4D attack. This evaluation focused on the more complex concepts $C_{\text{unsafe}}$ which pose additional challenges for safe generation. Our analysis revealed that SteerDiff marginally outperformed SLD MAX, maintaining an Attack Success Rate (ASR) of around 30%. This discrepancy may be attributable to the broader and more ambiguous scope of larger-scale unsafe concepts, which complicates effective defense, as also observed by (Ma et al., 2024). Nonetheless, SteerDiff remains the most competitive defense across all categories. Detailed quantitative results for each unsafe concept are presented in subsection A.3.
Finally, we assessed all safe T2I models on the NSFW-200 dataset using the SneakyPrompt attack methodology (Yang et al., 2024). SteerDiff achieving 7.5% ASR. In comparison, SLD STRONG and SLD MAX exhibited higher ASRs of 14.5% and 8.5%, respectively. These results underscore the effectiveness of SteerDiff in defending against sophisticated red-teaming attacks.
@@
4.4 RQ3: Can SteerDiff maintain high image fidelity and text alignment?
We have demonstrated the effectiveness of SteerDiff in mitigating generation of inappropriate content in T2I diffusion models. However, maintaining high image fidelity and ensuring strong alignment between generated images and input text prompts are equally important. Ideally, SteerDiff should have minimal or no impact on prompts that are already safe. To assess these aspects, we evaluated SteerDiff with COCO FID-30K score for image fidelity and CLIP score for measuring alignment between generated images and input text prompts.
As shown in Table 3, SteerDiff achieved a lower FID-30K score (15.45) compared to baseline models [^2] FID-30K score of SLD STRONG and SLD MAX are sourced from  Schramowski et al. (2023)., indicating better image fidelity. While CLIP score (0.78) is slightly lower than SD1.4 and ESD, it remains competitive among other methods, demonstrating that SteerDiff has minimal impact on text-image alignment specificity.
Table 3: SteerDiff shows better image fidelity and text alignment performance across all methods on COCO 30k images. All methods show good CLIP score consistency with SD.
@@ 4.4 RQ3: Can SteerDiff maintain high image fidelity and text alignment?
Method FID-30K $\downarrow$ CLIP $\uparrow$
SD1.4 16.47 0.81
@@
5 Discussion on Concept Forgetting

### v2

@@ 2.3. Steering Toward Safe Content
To guide the generation process toward safe outputs, we propose a transformation that maps unsafe prompt embeddings to safe regions while preserving their original semantic meaning. Our key insight is that linear transformations in the word embedding space can effectively alter the generation style, as demonstrated in prior work (Han et al., 2023). Since SteerDiff operates directly on these embeddings, this approach allows us to adjust unsafe representations before they are processed by the diffusion model.
Formally, let $e_{\text{unsafe}}$ denote the embedding of an unsafe phrase. We define the steered embedding as:
(3) $e_{\text{steered}}=\epsilon\cdot W\cdot e_{\text{unsafe}}+(1-\epsilon)\cdot e_{\text{unsafe}}$
where $W$ is a learnable transformation matrix and $\epsilon\in[0,1]$ is a hyperparameter controlling the intensity of the transformation. This formulation allows for fine-grained control, interpolating between the original and transformed embeddings.
To learn the transformation matrix $W$, we utilize a supervised learning framework on a paired dataset consisting of unsafe phrases and their corresponding safe counterparts, as described in subsection 2.1. The training objective minimizes the discrepancy between the transformed embedding and the target safe embedding:
(4) $\mathcal{L}=\left\|e_{\text{safe}}-W\cdot e_{\text{unsafe}}\right\|^{2}$
where $e_{\text{safe}}$ is the embedding of the corresponding safe phrase in the paired dataset. This supervised learning framework ensures that the transformation learns meaningful semantic shifts toward safety, enabling the system to retain user intent while suppressing unsafe visual outputs.
Although the transformation matrix $W$ is trained on curated phrase pairs, the steering mechanism is designed to be modular and efficiently updatable. As safety policies or blacklists evolve, new safe–unsafe pairs can be incrementally added without requiring full retraining from scratch. Due to the relatively low dimensionality and simplicity of the linear transformation, $W$ can be re-optimized rapidly using a small number of new examples (e.g., via fine-tuning or few-shot adaptation). This allows the system to dynamically accommodate new safety requirements with minimal engineering overhead, making it practical for deployment in fast-changing moderation contexts.
While our framework employs a unified transformation matrix $W$ to steer unsafe prompt embeddings, it is important to clarify that $W$ encapsulates a multidimensional safety subspace that naturally attenuates various unsafe concepts. In our approach, the sliding window mechanism identifies unsafe segments within a prompt, and the same transformation is applied uniformly. As a result, while the framework is effective in steering the overall unsafe content toward safe regions, it does not currently support dynamic, selective suppression of individual unsafe aspects (e.g., reducing violent content while retaining nudity-related elements). We acknowledge this as a limitation of our fixed-$\epsilon$ design and consider the development of concept-specific steering strategies as future work.
2.4. Inference
@@
Once SteerDiff is trained, it can be seamlessly integrated into diffusion models without additional fine-tuning. In practical applications, SteerDiff effectively detects the presence of blacklisted concepts and steers inappropriate prompts toward generating safe images.
@@
Consider a T2I model equipped with a text encoder. As illustrated in Figure 1, a user provides a prompt, which can be either safe or unsafe. The input prompt is first embedded by a text encoder. We define a concept blacklist $C_{\text{unsafe}}$ that contains potentially inappropriate concepts. Then the identifier detects whether prompt $p$ contains any inappropriate concepts $c^{\prime}\in C_{\text{unsafe}}$. If such a concept is detected, a linear transformation is applied to the prompt’s embedding to steer it toward safer content. Specifically, this transformation alters the latent representation to avoid generating inappropriate content without compromising the semantics of the original prompt. Once the transformation is applied, the modified embedding is passed through the diffusion model to generate a safe image. The transformed embedding ensures that the resulting image adheres to Safe-for-work guidelines while preserving the user’s original intent.
@@
Our method offers three key advantages: efficiency, effectiveness, and versatility. By operating at the prompt embedding level, it eliminates the need for computationally intensive model retraining while preserving the original semantics, allowing SteerDiff to act as an intermediary between user input and the diffusion model. This approach prevents unsafe content at the earliest stage of generation, blocking unsafe latent representations directly and reliably. Additionally, this steering process is lightweight and easily adaptable, enabling SteerDiff to align generated images with ethical and safety standards without introducing significant delays in the image generation pipeline. The transformation can be dynamically updated as blacklists or safety criteria evolve, making the system robust and scalable for various concept removal tasks in real-world applications.
3. Experiment
3.1. Benchmarks and Metrics
We evaluate our safety mitigation technique using multiple benchmarks that assess both the reduction of inappropriate content and the maintenance of image quality and text alignment. Lastly, we analyze SteerDiff’s versatility via an artist style removal task.
Inaproppriate Image Prompts (I2P)
The I2P benchmark (Schramowski et al., 2023) consists of 4,703 real user prompts likely to produce inappropriate images. We measure performance using two metrics: Inappropriateness Rate: The ratio of inappropriate images out of all images generated with the vanilla method. Attack Success Rate (ASR): The percentage of images flagged as inappropriate when adversarial red-teaming techniques are applied.
MS-COCO FID-30K
We adopt the MS-COCO FID-30K benchmark (Pavlov et al., 2023), which comprises 30,000 captions sampled from the MS-COCO validation set (Lin et al., 2014). This benchmark has become a standard for evaluating text-to-image models. Performance is quantified using the Fréchet Inception Distance (FID) and CLIP scores, which respectively measure image fidelity and the semantic alignment between generated images and input prompts.
Artist Style Removal
To demonstrate versatility, we evaluate our method on an artist style removal task following the setup from ESD (Gandikota et al., 2023). The artist removal dataset comprises five modern artists and thematic styles: Kelly McKernan, Thomas Kinkade, Tyler Edlin, Kilian Eng, and the series “Ajin: Demi-Human”. For each artist, 40 generic prompts are composed to invoke the artist’s style. We assess style removal using the Learned Perceptual Image Patch Similarity (LPIPS) metric (Zhang et al., 2018). Specifically, LPIPSe is calculated on the erased artist. A higher LPIPSe value indicates more effective style removal. LPIPSu is calculated on unerased artists, with a lower LPIPSu suggesting minimal impact on unrelated artists.
3.2. Models and Baseline
To evaluate the efficacy of our safety mitigation technique, we employ Stable Diffusion (SD) v1.4 as the base model. SD v1.4 is widely adopted in prior works on concept removal and red-teaming, making it the most relevant benchmark. For each prompt, three images are generated to reduce randomness, as generating only one image might coincidentally omit inappropriate content. We define a general unsafe concept set $C_{\text{unsafe}}$, applicable across all categories in I2P (details refer to supplementary materials).
@@ 3.2. Models and Baseline
We evaluated inappropriate image content by combining two classifiers. First, Q16 classifier (Schramowski et al., 2022), which is also used for annotating LAION-5B dataset (Schuhmann et al., 2022), was utilized to detect a broad range of inappropriate content. Second, we applied NudeNet (Praneeth, 2019) to identify sexually explicit content. For nude content, we classified exposed buttocks, exposed breasts, exposed genitalia, and exposed anus content as inappropriate, while allowing otherwise provocative images (Schramowski et al., 2023; Wu et al., 2024; Ma et al., 2024). Unless otherwise specified, an image was classified as inappropriate if at least one of the classifiers flagged it as containing such content.
Compare Methods
For the following RQ1, we evaluated SteerDiff with the state-of-the-art defense model SPM (Lyu et al., 2024), Erased Stable Diffusion (ESD) (Gandikota et al., 2023), MACE (Lu et al., 2024), RACE (Kim et al., 2024), AdvUnlearn (Zhang et al., 2024a) and Safe Latent Diffusion (SLD) (Schramowski et al., 2023) under its MAX and STRONG configurations. Note that SPM, ESD, MACE, RACE, and AdvUnlearn can only defend the generation of nude images in their implementation. For RQ2, we selected state-of-the-art red-teaming frameworks, including Prompting4Debugging (P4D)(Chin et al., 2023), SneakyPrompt(Yang et al., 2024b), Ring-A-Bell (Tsai et al., 2023), and MMA-Diffusion (Yang et al., 2024a). Following the developers’ standard instructions, we conducted our evaluation on the I2P dataset with the P4D testbed, Ring-A-Bell nude, MMA-nudity, and NSFW_200 dataset (Yang et al., 2024b), utilizing the standard configurations provided by each framework.
3.3. RQ1: Effectiveness of SteerDiff
@@
To investigate the ability of SteerDiff in both identifying and steering inappropriate concepts, we started by demonstrating its effectiveness in reducing the generation of explicit content. Next, we expanded the scope of inappropriate concepts to $C_{\text{unsafe}}$ to investigate whether SteerDiff could effectively identify and steer prompts containing a wider range of inappropriate content toward generating safe images. Notably, we used inappropriate terms from SteerDiff’s training set as the naive blacklist.
Table 1. SteerDiff effectively removes sexual content from SD v1.4 on the I2P dataset, outperforming defend methods ESD (Gandikota et al., 2023), SPM (Lyu et al., 2024), MACE (Lu et al., 2024), AdvUnlearn (Zhang et al., 2024a), SLD STRONG, and SLD MAX (Schramowski et al., 2023).
Method Inappropriate probability % $\downarrow$
I2P(Sexual) (Schramowski et al., 2023)
SD-v1.4 46.29
ESD (Gandikota et al., 2023) 37.52(-8.77)
SPM (Lyu et al., 2024) 23.39(-22.90)
SLD STRONG (Schramowski et al., 2023) 32.65(-13.64)
SLD MAX (Schramowski et al., 2023) 21.04(-25.25)
MACE (Lu et al., 2024) 17.43(-28.86)
RACE (Kim et al., 2024) 10.22(-36.07)
AdvUnlearn (Zhang et al., 2024a) 21.13(-25.16)
SteerDiff (Ours) 16.33(-29.96)
@@
Table 2. SteerDiff demonstrates the best performance in reducing the probability of generating inappropriate content (where lower values are better). The probabilities shown represent the likelihood of generating images classified as inappropriate by combining the Q16 and NudeNet classifiers across various I2P categories. The best performances are bolded, and the second-best performances are underscored.
@@
SD-v1.4 27.27 19.05 27.65 30.34 46.29 35.98 18.16 30.17
Blacklist 19.48(-7.79) 14.68(-4.37) 17.99(-9.66) 18.23(-12.11) 21.59(-24.70) 22.31(-13.67) 10.87(-7.29) 17.86
SLD STRONG 6.49(-20.78) 6.80(-12.25) 5.42(-22.23) 5.24(-25.10) 12.67(-33.62) 11.33(-24.65) 3.03(-15.13) 7.97
SLD MAX 3.90(-23.37) 3.76(-15.29) 5.29(-22.36) 2.25(-28.09) 8.38(-37.91) 6.31(-29.67) 2.75(-15.41) 5.15
SteerDiff (Ours) 5.63(-21.64) 4.25(-14.80) 2.91(-24.74) 4.74(-25.60) 6.89(-39.40) 6.78(-29.20) 3.99(-14.17) 4.51
Explicit Content Removal
We first evaluate the performance of SteerDiff against competitive methods on the I2P dataset, focusing on the sexual content category. Table 1 reports the inappropriate probability for each method, where the values in parentheses denote the percentage reduction with respect to the SD-v1.4 baseline. The baseline SD-v1.4 model exhibits an inappropriate probability of 46.29%. Post-hoc methods, such as ESD and SPM, reduce this probability to 37.52% and 23.39%, respectively. More advanced techniques like SLD MAX, MACE, and RACE further lower the probability. RACE achieves the best result with 10.22%. Meanwhile, our proposed SteerDiff obtains a value of 16.33%, representing a significant reduction (29.96%). These quantitative results demonstrate that SteerDiff effectively mitigates explicit content generation.
Inappropriate Content Removal
@@
We further investigated a more comprehensive inappropriate set $C_{\text{unsafe}}$ defined in Equation 1. We began our evaluation by demonstrating the inappropriate generation of SD v1.4 without any safety measures, as well as a basic blacklist-based prompt filtering approach. Table 2 presents the probability of generating inappropriate content for each category. Varying from different categories, SD v1.4 generated inappropriate content with probabilities ranging from 18.16% to 46.29%. The naive blacklist approach slightly reduced probability, but inappropriate content was still generated in 17.86% of cases across all categories. This confirms that simple blacklists, while providing minimal mitigation, remain ineffective as comprehensive safeguards due to their inability to address the semantic complexity and evolving nature of unsafe content.
@@
Next, we evaluated our proposed method SafeDiff against state-of-the-art safety mechanisms SLD STRONG and SLD MAX across various I2P concepts. As shown in Table 2, SafeDiff demonstrated superior performance, reducing inappropriate content generation by over 95%, with SLD MAX ranking second in effectiveness. In particular, SteerDiff outperformed its closest competitor, SLD MAX, in categories of violence, sexual content, and overall inappropriate content. As a result, only 5% of images generated by SteerDiff were still classified as inappropriate. However, it is worth noting that the Q16 and Nudenet classifiers tend to flag images as inappropriate even when problematic content has been significantly reduced. In summary, SteerDiff effectively mitigates the generation of inappropriate content through targeted identification and modification of unsafe concepts within the text embedding space, offering a more robust solution compared to existing approaches.
3.4. RQ2: Effectivenss of SteerDiff under red-teaming attacks
SafeDiff achieves state-of-the-art performance on the I2P dataset; robustness against red-teaming attacks remains a critical criterion for effective safety defense. Red-teaming strategies vary significantly in nature. Black-box attacks typically paraphrase or substitute unsafe content with semantically similar alternatives to evade detection, whereas white-box attacks leverage model gradients or internal representations to bypass safety mechanisms. To comprehensively assess robustness, we first evaluate SafeDiff under the black-box red-teaming attack P4D (Chin et al., 2023). Subsequently, we analyze its performance against three representative white-box attacks: SneakyPrompt (Yang et al., 2024b), Ring-A-Bell (Tsai et al., 2023), and MMA-Diffusion (Yang et al., 2024a).
Table 3. Attack success rate (% $\downarrow$) on I2P and P4D benchmarks.
@@ 3.4. RQ2: Effectivenss of SteerDiff under red-teaming attacks
I2P (all) P4D (nude) P4D (all)
SD-v1.4 30.17 95.21 98.70
ESD - 55.40(-39.81) -
SLD STRONG 7.97(-22.20) 48.21(-47.00) -
SLD MAX 5.15(-25.02) 37.25(-57.96) 30.95(-67.75)
SteerDiff (Ours) 4.51(-25.66) 25.36(-69.85) 29.16(-69.54)
White-box Attack
Table 3 To investigate the effectiveness of concept removal approaches, we first focused on the “nudity” category, as it is commonly recognized as explicitly harmful in the context of generative models. Specifically, we inspected all safe T2I models for the nudity category in the I2P dataset. As shown in LABEL:T:exp2, ESD, SPM, and SLD exhibited poor performance when countering nudity-related attacks. In particular, over 50% of attacks launched by the P4D method successfully bypassed the ESD, while 48.21% and 37.25% circumvented SLD STRONG and MAX defenses, respectively. In contrast, SteerDiff achieves a significantly lower attack success rate of 25.36%, marking it as the most effective defense against red-teaming attacks targeting nudity.
Next, we evaluated the robustness of SLD MAX and SteerDiff across all categories within the I2P dataset under the P4D attack, excluding ESD and SLD STRONG due to limited category support. Results in LABEL:T:exp2 revealed that SteerDiff marginally outperformed SLD MAX, maintaining an Attack Success Rate of around 30%. This discrepancy may be attributable to the broader and more ambiguous scope of larger-scale unsafe concepts, which complicates effective defense, as also observed by (Ma et al., 2024). Nonetheless, SteerDiff remains the most competitive defense across all categories. Detailed quantitative results for each unsafe concept and image results are presented in the supplementary materials.
Table 4. Attack success rate (% $\downarrow$) under SneakyPrompt, Ring-A-Bell, and MMA-Diffusion.
Method Attack success rate (ASR) % $\downarrow$
SneakyPrompt Ring-A-Bell MMA-Diffusion
SD-v1.4 69.35 83.10 95.70
ESD 51.00(-18.35) 52.80(-30.30) 87.30(-8.40)
SPM 34.00(-35.35) 34.18(-48.92) 73.30(-22.40)
SLD STRONG 14.50(-54.85) 62.00(-21.10) 92.00(-3.70)
SLD MAX 8.50(-60.85) 57.50(-25.60) 83.70(-12.00)
SteerDiff (Ours) 7.50(-61.85) 27.85(-55.25) 28.80(-66.90)
Black-box Attack
Lastly, we assessed all safe T2I models under SneakyPrompt, Ring-A-Bell, and MMA-Diffusion. As shown in Table 4, SteerDiff achieving 7.5%, 27.85%, and 28.80% ASR, respectively. In comparison, ESD, SPM, SLD STRONG, and SLD MAX exhibited higher ASRs across all attack types. These results underscore the effectiveness of SteerDiff in defending against sophisticated red-teaming attacks.
@@
Table 5. Image fidelity and text alignment on COCO 30K.
Method MS-COCO FID-30K
@@
FID-30K $\downarrow$ CLIP $\uparrow$
SD-v1.4 - 0.81
@@
3.5. RQ3: Generated Images Quality
We have demonstrated the effectiveness of SteerDiff in mitigating the generation of inappropriate content in T2I diffusion models. However, maintaining high image fidelity and ensuring strong alignment between generated images and input text prompts are equally important. Ideally, SteerDiff should have minimal or no impact on prompts that are already safe. To assess these aspects, we evaluated SteerDiff with MS-COCO FID-30K score for image fidelity and CLIP score for measuring alignment between generated images and input text prompts.
As shown in Table 5, SteerDiff achieves a FID-30K score of 15.45—the second-lowest among all baselines , indicating better image fidelity. While CLIP score (0.78) is slightly lower than ESD and SPM, it remains competitive among other methods, demonstrating that SteerDiff has minimal impact on text-image alignment specificity.
3.6. Artist Style Removal
In this section, we explore the potential of SteerDiff to erase specific artist styles. For this experiment, we follow the artist removal dataset from ESD (Gandikota et al., 2023), where SteerDiff was applied to remove references to “Van Gogh” and “Kelly McKernan”, while maintaining other artist styles, such as “Pablo Picasso”, “Thomas Kinkade”, “Tyler Edlin”, and “Kilian Eng”.